# Accelerating Volumetric Medical Image Annotation via Short-Long Memory SAM 2

### v1

@@ Abstract
Manual annotation of volumetric medical images, such as magnetic resonance imaging (MRI) and computed tomography (CT), is a labor-intensive and time-consuming process. Recent advancements in foundation models for video object segmentation, such as Segment Anything Model 2 (SAM 2), offer a potential opportunity to significantly speed up the annotation process by manually annotating one or a few slices and then propagating target masks across the entire volume. However, the performance of SAM 2 in this context varies. Our experiments show that relying on a single memory bank and attention module is prone to error propagation, particularly at boundary regions where the target is present in the previous slice but absent in the current one. To address this problem, we propose Short-Long Memory SAM 2 (SLM-SAM 2), a novel architecture that integrates distinct short-term and long-term memory banks with separate attention modules to improve segmentation accuracy. We evaluate SLM-SAM 2 on three public datasets covering organs, bones, and muscles across MRI and CT modalities. We show that the proposed method markedly outperforms the default SAM 2, achieving average Dice Similarity Coefficient improvement of 0.14 and 0.11 in the scenarios when 5 volumes and 1 volume are available for the initial adaptation, respectively. SLM-SAM 2 also exhibits stronger resistance to over-propagation, making a notable step toward more accurate automated annotation of medical images for segmentation model development.
@@ 1 Introduction
Semantic segmentation plays a crucial role in medical image analysis, serving as a foundational step in numerous clinical and research applications, including tumor identification and measurement, treatment planning, and body composition analysis [1, 2, 3]. With advances in computer vision, deep learning-based methods (e.g., UNet [4]) have increasingly become the dominant approach for medical image segmentation due to their superior robustness and efficiency. However, developing reliable deep learning models often demands high-quality manual annotations, which are time-consuming and expensive to acquire [5], particularly for 3-dimensional (3D) volumetric medical data (e.g., MRI and CT) that consist of numerous 2D slices, requiring detailed labeling.
3D medical imaging data typically consists of a stack of 2D slices acquired along an additional spatial axis. These 2D consecutive slices are often highly correlated, capturing gradual anatomical changes across the axis. Consequently, the 3D volume data can be interpreted as a temporal progression, analogous to frames in a video [6, 7]. Leveraging such similarity between adjacent slices offers an effective method to accelerate annotation process by propagating annotations from a single slice to the rest of the volume [8, 9, 10, 11, 6, 7].
@@
The recently developed video object segmentation (VOS) foundation model, Segment Anything Model 2 (SAM 2) [12], has been applied to medical imaging segmentation [6], offering an alternative. Vision foundation models have demonstrated exceptional generalization capabilities without requiring access to large-scale labeled datasets [13, 14, 2]. Through extensive pretraining, these models acquire robust and adaptable feature representations, which facilitate their applications to downstream tasks with limited annotated data [2]. In VOS tasks, the objective is to track objects throughout a video sequence by propagating an initial object mask provided in the first frame [15, 12]. By interpreting the depth dimension as the temporal axis, SAM 2 offers a supervised method (i.e., training with images and corresponding masks) to propagate annotation masks across slices. For instance, SAM 2 has been utilized to build a 3D interaction framework [7], and an extension based on SAM 2 has been developed and integrated into 3D Slicer to facilitate the annotation of 3D volumes [16].
Based on SAM [13], SAM 2 enables video object segmentation by introducing a memory encoder, memory bank, and memory attention [12]. The memory encoder constructs memory representation through element-wise summation of encoded images and their corresponding output masks. The memory bank serves as a repository that retains information and past predictions from $N$ recent slices and $M$ prompted slices. Additionally, the memory attention module enhances the feature representation of the current slice by conditioning it on the features and predictions of past slices from the memory bank. The output features from the memory attention are then input to the mask decoder to generate mask prediction.
@@
Intuitively, a longer memory (i.e., larger $N$) allows the model to track gradual changes in the object, thereby enhancing the stability and continuity of predictions over slices. In contrast, a shorter memory (i.e., smaller $N$) emphasizes the most recent slices, improving the model’s adaptability to discontinuous changes in object appearance. To leverage the complementary advantages of both long-term and short-term memory, we propose Short-Long Memory SAM 2 (SLM-SAM 2)  [^1] GitHub at: https://github.com/mazurowski-lab/SLM-SAM2, a novel framework that incorporates two memory banks, one long-term and one short-term, to achieve more robust and accurate segmentation. We validate the effectiveness of SLM-SAM 2 through experiments on three public datasets spanning both MRI and CT modalities. The results demonstrate that SLM-SAM 2 significantly outperforms existing 2D and 3D segmentation methods. Our contributions are summarized below:
@@
  • Our method significantly outperforms existing approaches in segmenting organs, bones, and muscles across both MRI and CT modalities, achieving an average dice score improvement of at least $7\%$ across different datasets.
  • We demonstrate both qualitatively and quantitatively that SLM-SAM 2 can effectively alleviate the over-propagation issue.
@@
Figure 1: Overview of SLM-SAM 2. (a) is the general pipeline of SLM-SAM 2. The dynamic short-long memory module consists of short-term and long-term memory banks ($\mathcal{M}_{short}$ and $\mathcal{M}_{long}$), two distinct attention modules for each of them ($\mathcal{A}_{short}$ and $\mathcal{A}_{long}$), and the dynamic attention fuser ($\mathcal{F}$). Specifically, $\mathcal{M}_{long}$ includes one annotated slice and up to six additional slices, while $\mathcal{M}_{short}$ contains only the most recent slice. (b) illustrates the architecture of the dynamic attention fuser $\mathcal{F}$. (c) provides an explanation of the key notations used in SLM-SAM 2.
@@ 2.1 Semantic Segmentation in Medical Imaging
Semantic segmentation aims to classify each pixel in an image, making it an essential task in both natural and medical imaging [17, 18]. Convolutional Neural Networks (CNNs) have been a dominant approach for vision tasks, with UNet [4] emerging as a widely adopted model in medical image segmentation. The skip connections between encoder and decoder allows UNet to effectively capture both global and local context. Building on UNet, nnUNet [19] introduces a fully automated training pipeline configuration, which can be efficiently adapted to any new medical segmentation tasks. Nevertheless, one of the key limitations of CNNs is that they often struggle to capture long-range dependencies [20].
To overcome this issue, Vision Transformers (ViTs) employ a transformer architecture that divides images into sequence of patches [21]. The self-attention mechanism in ViTs allows them to capture global context, enhancing segmentation accuracy for anatomies with large receptive fields (e.g., the liver). Recently, vision foundation models, such as SAM [16], have further advanced the medical image segmentation. SAM leverages a transformer-based encoder and extensive pretraining on large-scale datasets, enabling exceptional generalization capabilities across multiple downstream tasks with minimal labeled data [22, 23].
@@ 2.2 Video Object Segmentation (VOS)
Video object segmentation focuses on segmenting particular object instance given an entire video sequence and an object mask on the first frame, either manual annotated or automatically generated [15]. VOS has numerous applications in various domains, including autonomous vehicles [24, 25], robotics [26, 27] and medical imaging analysis [6, 28]. VOS methods can be broadly categorized into unsupervised VOS and semi-supervised VOS [29]. Unsupervised VOS aims to identify and segment the most salient objects in a video without any given annotations. In contrast, semi-supervised VOS requires initial annotations, which are typically provided in the first frame [30].
Recently, several novel approaches, such as Cutie [31] and SAM 2 [12], have been introduced for semi-supervised VOS tasks. Cutie employs an iterative interaction between top-down object-level memory reading and bottom-up pixel-level features to effectively capture the semantic structure of foreground objects [31]. SAM 2 incorporates a pretrained ViT backbone along with a memory mechanism that includes a memory bank, memory encoder, and memory attention module [12]. The enriched feature representations from the pretrained ViT encoder allow SAM 2 to achieve robust generalization performance across different tasks.
@@ 2.3 Medical Image Annotation for Segmentation
To address these challenges, various approaches have been developed, which can be broadly categorized into unsupervised, supervised and interactive methods. Unsupervised methods focus on learning spatial relationship between adjacent slices in 3D data, enabling label propagation from a single annotated slice to entire volume during inference [8, 11]. However, these methods are prone to error drift [11]. Supervised methods often utilize few-shot learning [35, 36], where a model is trained on a small, fully annotated subset data before being applied to the remaining data. This category involves 2D [37, 38], 3D [39, 40], and VOS models [31, 12], with VOS technique sometimes requiring minimal annotations during the inference stage. Interactive annotation methods integrate automated predictions with real-time human corrections, enabling quick refinement of initial annotations, and thus improving annotation efficiency [35, 41].
@@ 3 Methods
This work aims to develop robust and reliable slice propagation models that can accelerate annotation of volumetric medical image data. In this section, we define the problem and introduce the proposed SLM-SAM 2 in detail.
@@ 3.1 Problem Definition
Given a fully annotated volume set $\{X_{i}|i=0,...,n\}$ for training, the objective is to segment any incoming volumes from $\{Y_{i}|i=0,...,m\}$ by annotating a single slice from each volume. In this study we chose $n=5$, reflecting a practical setting in which users initially provide full annotations for a limited number of volumes and subsequently annotate only one slice for each newly acquired volume.
@@ 3.2 Model Architecture of SLM-SAM 2
Fig. 1 (a) depicts the overall SLM-SAM 2 pipeline. SLM-SAM 2 retains the foundational architecture of SAM 2, including the image encoder $E_{I}$, prompt encoder $E_{P}$, and mask decoder $D$. In SAM 2, at slice $s_{i}$, the memory attention $A_{SAM2}$ can be simply expressed as:
$A_{SAM2}=\mathcal{A}(E_{I}(s_{i}),[E_{M}(s_{j})|s_{j}\in\mathcal{M}])$ (1)
where $E_{M}$ is the memory encoder and $\mathcal{M}$ is the memory bank. We use the SAM 2 “tiny” model architecture in this work.
The key dynamic short-long memory module consists of a short-long memory bank mechanism and a dynamic attention fuser, as described below.
3.2.1 Short-long Memory Bank
@@
$N$: Number of training epochs.
@@
1 for $\text{epoch}\leftarrow 1$ to $N$ do
@@ 4            $\boldsymbol{F}_{P}\leftarrow E_{P}(P)$;
5             $L_{\text{total}}\leftarrow 0$;
@@
                   // Forward pass
7                   $\boldsymbol{F}_{I}\leftarrow E_{I}(I_{t})$;
8                   $A_{long}\leftarrow\mathcal{A}_{long}(E_{I},\mathcal{M}_{long})$;
9                   $A_{short}\leftarrow\mathcal{A}_{short}(E_{I},\mathcal{M}_{short})$;
10                   $A_{SLM-SAM2}\leftarrow\mathcal{F}(A_{long},A_{short})$;
11                   $(S_{t}^{pred},Q_{t}^{pred})\leftarrow D(A_{SLM-SAM2},\boldsymbol{F}_{P})$;
@@ 14                   $L_{\text{total}}\leftarrow L_{\text{total}}+L_{\text{step}}$;
16                   $\boldsymbol{F}_{M}\leftarrow E_{M}(I_{t},S_{t}^{pred})$;
17                   $\mathcal{M}_{long}\leftarrow\textsc{Update}(\mathcal{M}_{long},E_{m})$;
18                   $\mathcal{M}_{short}\leftarrow\textsc{Update}(\mathcal{M}_{short},E_{m})$;
19
@@ 22             Optimizer_Step();
Similar to SAM 2, SLM-SAM 2 inputs the new incoming slices to the image encoder to generate the corresponding feature representations. To enable both short-term and long-term memory storage, SLM-SAM 2 introduces an additional memory bank and memory attention. The long-term memory bank $\mathcal{M}_{long}$ follows the original SAM 2 design by storing $N_{long}$ recent and $M_{long}$ prompted slices, ensuring continuity in segmentation. The short-term memory bank $\mathcal{M}_{short}$ sorely retains the most recent $N_{short}$ slice, where $N_{short}$ is a typically small positive integer, allowing model predictions to quickly adapt to abrupt object changes. During propagation, the conditions of the current slices on both short-term and long-term memory, $A_{short}$ and $A_{long}$, are computed independently through two memory attention modules, $\mathcal{A}_{short}$ and $\mathcal{A}_{long}$, respectively.
$A_{short}=\mathcal{A}_{short}(E_{I}(s_{i}),[E_{M}(s_{j})|s_{j}\in\mathcal{M}_{short}])$ (2)
$A_{long}=\mathcal{A}_{long}(E_{I}(s_{i}),[E_{M}(s_{j})|s_{j}\in\mathcal{M}_{long}])$ (3)
The two memory modules are of the same architecture, resulting in output attentions of the same size. Specfically, SLM-SAM 2 uses $N_{long}=6$, $M_{long}=1$ and $N_{short}=1$.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Semantic segmentation plays a crucial role in medical image analysis, serving as a foundational step in numerous clinical and research applications, including tumor identification and measurement, treatment planning, and body composition analysis [1, 2, 3, 4]. With advances in computer vision, deep learning-based methods (e.g., UNet [5]) have increasingly become the dominant approach for medical image segmentation due to their superior robustness and efficiency. However, developing reliable deep learning models often demands high-quality manual annotations, which are time-consuming and expensive to acquire [6], particularly for 3-dimensional (3D) volumetric medical data (e.g., MRI and CT) and medical videos (e.g., ultrasound videos) that consist of numerous 2D slices, requiring detailed labeling.
3D medical imaging data typically consists of a stack of 2D slices acquired along an additional spatial axis. These 2D consecutive slices are often highly correlated, capturing gradual anatomical changes across the axis. Consequently, the 3D volume data can be interpreted as a temporal progression, analogous to frames in a video [7]. Leveraging such similarity between adjacent slices offers an effective method to accelerate the annotation process by propagating annotations from a single slice to the rest of the volume [8, 9, 10, 11].
@@
The recently developed video object segmentation (VOS) foundation model, Segment Anything Model 2 (SAM 2) [12], has been applied to medical imaging segmentation [13, 7], offering an alternative. Vision foundation models have demonstrated exceptional generalization capabilities without requiring access to large-scale labeled datasets [14, 15, 4]. Through extensive pretraining, these models acquire robust and adaptable feature representations, which facilitate their applications to downstream tasks with limited annotated data [4]. In VOS tasks, the objective is to track objects throughout a video sequence by propagating an initial object mask provided in the first frame [16, 12]. By interpreting the depth dimension as the temporal axis, SAM 2 offers a supervised method (i.e., training with images and corresponding masks) to propagate annotation masks across slices. For instance, SAM 2 has been utilized to build a 3D interaction framework [7].
Based on SAM [14], SAM 2 enables video object segmentation by introducing a memory encoder, memory bank, and memory attention [12]. The memory encoder constructs memory representation through element-wise summation of encoded images and their corresponding output masks. The memory bank serves as a repository that retains information and past predictions from $\boldsymbol{N}$ recent slices and $\boldsymbol{M}$ prompted slices. Additionally, the memory attention module enhances the feature representation of the current slice by conditioning it on the features and predictions of past slices from the memory bank. The output features from the memory attention are then input to the mask decoder to generate a mask prediction.
@@
Intuitively, a longer memory (i.e., larger $N$) allows the model to track gradual changes in the object, thereby enhancing the stability and continuity of predictions over slices. In contrast, a shorter memory (i.e., smaller $N$) emphasizes the most recent slices, improving the model’s adaptability to discontinuous changes in object appearance. To leverage the complementary advantages of both long-term and short-term memory, we propose Short-Long Memory SAM 2 (SLM-SAM 2)  [^1] GitHub at: https://github.com/mazurowski-lab/SLM-SAM2, a novel framework that incorporates two memory banks, one long-term and one short-term, to achieve more robust and accurate segmentation. We validate the effectiveness of SLM-SAM 2 through experiments on four public datasets spanning MRI, CT, and ultrasound video modalities. The results demonstrate that SLM-SAM 2 significantly outperforms existing 2D and 3D segmentation methods. Our contributions are summarized below:
@@
  • Our method significantly outperforms existing approaches in segmenting organs, bones, and muscles across MRI, CT, and ultrasound videos, achieving an average dice score improvement of at least $7\%$ across different datasets.
  • We demonstrate both qualitatively and quantitatively that SLM-SAM 2 can effectively alleviate the over-propagation issue and other common propagation errors.
  • Our method significantly reduces the manual efforts in correcting propagated masks by $60.575\%$ in seconds per volume and $53.574\%$ in corrected-slice ratio compared to SAM 2.
@@ 2.1 Semantic Segmentation in Medical Imaging
Semantic segmentation aims to classify each pixel in an image, making it an essential task in both natural and medical imaging [17, 18]. Convolutional Neural Networks (CNNs) have been a dominant approach for vision tasks, with UNet [5] emerging as a widely adopted model in medical image segmentation. The skip connections between the encoder and decoder allow UNet to effectively capture both global and local context. Building on UNet, nnUNet [19] introduces a fully automated training pipeline configuration, which can be efficiently adapted to any new medical segmentation tasks. Nevertheless, one of the key limitations of CNNs is that they often struggle to capture long-range dependencies [20].
To overcome this issue, Vision Transformers (ViTs) employ a transformer architecture that divides images into a sequence of patches [21]. The self-attention mechanism in ViTs allows them to capture global context, enhancing segmentation accuracy for anatomies with large receptive fields (e.g., the liver). Recently, vision foundation models, such as SAM [22], have further advanced the medical image segmentation. SAM leverages a transformer-based encoder and extensive pretraining on large-scale datasets, enabling exceptional generalization capabilities across multiple downstream tasks with minimal labeled data [23, 13].
Recurrent networks, such as convolutional Long Short-Term Memory (ConvLSTM) and its variants, have also been explored in medical image segmentation [24, 25, 26, 27]. For instance, [26] replaced skip connection of UNet with Bidirectional ConvLSTM [28] to enhance intra-slice feature fusion. More related, [25] combined Bidirectional ConvLSTM with UNet to process full volumes segmentation sequentially, capturing dependencies across slices. Similarly, [27] introduced an LSTM multi-modal UNet to enforce inter-slice continuity for brain tumor segmentation. Nevertheless, these approaches rely on implicit hidden-state propagation to model inter-slice context and often lack object-specific memory mechanisms, limiting their capacity for fine-grained segmentation and robustness over long sequences.
Figure 1: Overview of SLM-SAM 2. (a) is the general pipeline of SLM-SAM 2. The dynamic short-long memory module consists of short-term and long-term memory banks ($\mathcal{M}_{short}$ and $\mathcal{M}_{long}$), two distinct attention modules for each of them ($\mathcal{A}_{short}$ and $\mathcal{A}_{long}$), and the dynamic attention fuser ($\mathcal{F}$). Specifically, $\mathcal{M}_{long}$ includes memories from one annotated slice and up to six additional slices, while $\mathcal{M}_{short}$ contains memory only from the most recent slice. $A_{short}$ and $A_{long}$ denote the outputs of $\mathcal{A}_{short}$ and $\mathcal{A}_{long}$, respectively, and $A_{SLM-SAM2}$ represents the fused attention features produced by $\mathcal{F}$. (b) illustrates the architecture of the dynamic attention fuser $\mathcal{F}$. (c) provides an explanation of the key notations used in SLM-SAM 2.
@@ 2.2 Video Object Segmentation (VOS)
Video object segmentation focuses on segmenting particular object instance given an entire video sequence and an object mask on the first frame, either manually annotated or automatically generated [16]. VOS has numerous applications in various domains, including autonomous vehicles [29, 30], robotics [31, 32], and medical imaging analysis [7, 33]. VOS methods can be broadly categorized into unsupervised VOS and semi-supervised VOS [34]. Unsupervised VOS aims to identify and segment the most salient objects in a video without any given annotations. In contrast, semi-supervised VOS requires initial annotations, which are typically provided in the first frame [35]. In addition, semi-supervised VOS methods also can be adopted for interactive VOS segmentation by combining with SAM [36, 37] or other 2D interactive segmentation methods [38].
Memory-based architectures have been widely adopted in advancing semi-supervised VOS. For example, STM [39] introduces a space-time memory network that stores past image-mask pairs in a single memory bank, retrieving them via a key-value attention mechanism. XMem [40] extends this paradigm with an Atkinson-Shiffrin-inspired design that maintains multiple memory stores for both long-term and short-term information, and consolidates frequently used prototypes based on cumulative affinity scores. Recently, additional approaches, such as Cutie [41] and SAM 2 [12], have been introduced for semi-supervised VOS tasks. Cutie employs an iterative interaction between top-down object-level memory reading and bottom-up pixel-level features to effectively capture the semantic structure of foreground objects [41]. SAM 2 incorporates a pretrained ViT backbone along with a memory mechanism that includes a memory bank, memory encoder, and memory attention module [12]. The enriched feature representations from the pretrained ViT encoder allow SAM 2 to achieve robust generalization performance across different tasks. Different from previous interactive VOS approaches, SAM 2 offers a unified model that directly accepts user prompts for real-time mask refinement in videos. Inspired by these prior works, our method builds on SAM 2 by integrating dual short-term and long-term memory banks and introducing a learnable fuser that dynamically combines the two memories. This design provides a novel mechanism to leverage SAM 2’s rich encoder and flexible prompting while enabling adaptive use of memory information for efficient volumetric medical image annotation.
@@ 2.3 Medical Image Annotation for Segmentation
To address these challenges, various approaches have been developed, which can be broadly categorized into unsupervised, supervised, and interactive methods. Unsupervised methods focus on learning spatial relationship between adjacent slices in 3D data, enabling label propagation from a single annotated slice to the entire volume during inference [8, 11]. However, these methods are prone to error drift [11]. Supervised methods often utilize few-shot learning [44, 45], where a model is trained on a small, fully annotated subset of data before being applied to the remaining data. This category involves 2D [46, 22], 3D [47, 48], and VOS models [41, 12], with VOS technique sometimes requiring minimal annotations during the inference stage. Interactive annotation methods integrate automated predictions with real-time human corrections, enabling quick refinement of initial annotations, and thus improving annotation efficiency [44].
@@ 3 Methods
This work aims to develop robust and reliable slice propagation models that can accelerate the annotation of volumetric medical image data. In this section, we define the problem and introduce the proposed SLM-SAM 2 in detail.
@@ 3.1 Problem Definition
Given a fully annotated volume set $\{X_{i}|i=0,...,n\}$ for training, the objective is to segment any incoming volumes from $\{Y_{i}|i=0,...,m\}$ by annotating a single slice from each volume. In this study, we chose $n=5$, reflecting a practical setting in which users initially provide full annotations for a limited number of volumes and subsequently annotate only one slice for each newly acquired volume.
@@ 3.2 Model Architecture of SLM-SAM 2
Fig. 1 (a) depicts the overall SLM-SAM 2 pipeline. SLM-SAM 2 retains the foundational architecture of SAM 2, including the image encoder $E_{I}$, prompt encoder $E_{P}$, and mask decoder $Dec$, but replaces its single memory bank and attention module with a novel dynamic short-long memory module, as shown in Fig. 1.
Given an input sequence of frames, the image encoder of SLM-SAM 2 produces a feature embedding of each frame. The initial prompt (e.g., manual annotations) is encoded by the prompt encoder. For mask prediction, each feature embedding is dynamically conditioned on short-term and long-term memories via separate learnable attention modules and then fused by a dynamic, trainable fusion module. The resulting embeddings are passed to the mask decoder to generate segmentation mask. The memory of each frame is generated from the memory encoder using its embedding features from the image encoder together with the corresponding predicted mask.
SLM-SAM 2 adopts the “tiny” Hiera image encoder [49] as the backbone, which is pretrained using MAE [50]. The prompt encoder of SLM-SAM 2 follows the same architecture of SAM 2 to allow mask inputs, and SLM-SAM 2 employs the same mask decoder architecture as SAM 2. The key dynamic short-long memory module consists of a short-long memory bank mechanism and a dynamic attention fuser, as described below.
3.2.1 Short-long Memory Bank Mechanism
@@
NN: Number of training epochs.
$\mathcal{M}_{long}$: Long-term memory bank
$\mathcal{M}_{short}$: Short-term memory bank
$\mathcal{A}_{long}$: Long-term memory attention module
$\mathcal{A}_{short}$: Short-term memory attention module
$\mathcal{F}$: Dynamic attention fuser
$E_{P},\boldsymbol{F}_{P}$: Prompt encoder and its output
$E_{I},\boldsymbol{F}_{I}$: Image encoder and its output
$E_{M},\boldsymbol{F}_{M}$: Memory encoder and its output
DecDec: Mask decoder
$S_{t}^{pred}$: Predicted mask at slice $t$
@@
1
21exfor $\text{epoch}\leftarrow 1$ to NN do
@@ 4    $\boldsymbol{F}_{P}\leftarrow E_{P}(P)$;
5    $(\boldsymbol{F}_{M})_{0}\leftarrow\boldsymbol{0}$;
    $L_{\text{total}}\leftarrow 0$ ;
    // total loss
@@ 5    $(\boldsymbol{F}_{M})_{0}\leftarrow\boldsymbol{0}$;
7
        1ex// Forward pass
8       $(\boldsymbol{F}_{I})_{t}\leftarrow E_{I}(I_{t})$;
9       $(A_{long})_{t}\leftarrow\mathcal{A}_{long}((\boldsymbol{F}_{I})_{t},\mathcal{M}_{long})$;
10       $(A_{short})_{t}\leftarrow\mathcal{A}_{short}((\boldsymbol{F}_{I})_{t},\mathcal{M}_{short})$;
11       $(A_{SLM-SAM2})_{t}\leftarrow\mathcal{F}((A_{long})_{t},(A_{short})_{t})$;
@@ 18       $(\boldsymbol{F}_{M})_{t}\leftarrow E_{M}(I_{t},S_{t}^{pred})$;
20       $\mathcal{M}_{long}\leftarrow\textsc{Update}(\mathcal{M}_{long},(\boldsymbol{F}_{M})_{t})$;
21       $\mathcal{M}_{short}\leftarrow\textsc{Update}(\mathcal{M}_{short},(\boldsymbol{F}_{M})_{t})$;
22
23
    // Backpropagate per sequence
@@ 26    Optimizer_Step();
In SAM 2, at slice $I_{t}$, the memory attention $A_{SAM2}$ can be expressed as  [^2] We denote slice, embedding, and memory at frame $t$ using $(.)_{t}$:
$(A_{SAM2})_{t}=\mathcal{A}((\boldsymbol{F}_{I})_{t},\mathcal{M})$ (1)
$(\boldsymbol{F}_{I})_{t}=E_{I}(I_{t})$ (2)
$\mathcal{M}=\{(\boldsymbol{F}_{M})_{j}|(\boldsymbol{F}_{M})_{j}\in\mathcal{M}\}$ (3)
where $(\boldsymbol{F}_{I})_{t}$ is the output embedding feature from the image encoder at slice $I_{t}$, $(\boldsymbol{F}_{M})_{j}$ represents memory at slice $I_{j}$, and $\mathcal{M}$ is the memory bank which stores these memories. The memory encoder uses a convolutional module to downsample the predicted mask and sum it element-wise with the corresponding image embedding, then applies a lightweight convolutional layer to produce the memory.
Similar to SAM 2, SLM-SAM 2 inputs the new incoming slices to the image encoder to generate the corresponding feature representations. To enable both short-term and long-term memory storage, SLM-SAM 2 introduces an additional memory bank and memory attention. The long-term memory bank $\mathcal{M}_{long}$ follows the original SAM 2 design by storing $N_{long}$ recent and $M_{long}$ prompted slices, ensuring continuity in segmentation. The short-term memory bank $\mathcal{M}_{short}$ solely retains the most recent $N_{short}$ slice, where $N_{short}$ is a typically small positive integer, allowing model predictions to quickly adapt to abrupt object changes. During propagation, the current slice’s features attend independently to the short-term memory and the long-term memory, producing attention outputs $A_{short}$ and $A_{long}$.
$(A_{short})_{t}=\mathcal{A}_{short}((\boldsymbol{F}_{I})_{t},\mathcal{M}_{short})$ (4)
$(A_{long})_{t}=\mathcal{A}_{long}((\boldsymbol{F}_{I})_{t},\mathcal{M}_{long})$ (5)
The two memory attention modules are of the same architecture, resulting in output attentions of the same size. Each module consists of a stack of $L$ transformer blocks. Specifically, SLM-SAM 2 uses $N_{long}=6$, $M_{long}=1$ and $N_{short}=1$. Consequently, the contents of the two memory banks at slice $i$ can be described as:
$\mathcal{M}_{short}=\{(\boldsymbol{F}_{M})_{t-1}\}$ (6)
$\mathcal{M}_{long}=\{(\boldsymbol{F}_{M})_{t-1},(\boldsymbol{F}_{M})_{t-2},...,(\boldsymbol{F}_{M})_{t-6},(\boldsymbol{F}_{M})_{c}\}$ (7)
where $(\boldsymbol{F}_{M})_{c}$ is the memory on the conditional slice (i.e., the initial annotated slice).
Each memory bank operates as a FIFO queue of recent-slice memories. At each update (in Algorithm 1), the short-term memory bank, $\mathcal{M}_{short}$, discards its current memory and enqueues the newly generated one. The long-term memory bank, $\mathcal{M}_{long}$, consistently retains the memory on conditional slices by appending the new memory but only dequeues its oldest entry on non-conditional slices.